In three more years Michael’s grandmother will be six times as old as Michael was last year when Michael’s present age is added to his grandfather’s present age the total is 68 how old is each one now
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

In three more years Michael’s grandmother will be six times as old as Michael was last year
g + 3 = 6(m - 1)
g + 3 = 6m - 6
g = 6m - 6 - 3
g = 6m - 9
:
when Michael’s present age is added to his grandmother’s present age the total is 68
m + g = 68
replace g with (6m-9)
m + 6m - 9 = 68
7m = 68 + 9
7m = 77
m = 11 yrs is M's age
then
g = 6(11) - 9
g = 66 - 9
g = 57 yrs is G's age
:
You should confirm this in both statements

Answer by ikleyn(52818)

.
In three more years Michael’s grandmother will be six times as old as Michael was last year when Michael’s present age
is added to his grandfather’s present age the total is 68 how old is each one now
~~~~~~~~~~~~~~~~~~~~~~~~


If you read the condition attentively,  you,  probably,  will notice that there are  3  (three,  THREE)  persons in the problem:

        - Michael, his grandmother and his grandfather.

Hence,  there are  3  (three,  THREE)  unknowns.

But the condition offers only  2  (two,  TWO)  conditions.


Instead of the conclusion:   There are too many unknowns and too few conditions in order for the problem could be solved properly.